what is the answer to 2(15 - 7g) = 3[ g + 2 - 3(2g - 1)]
Answer by gonzo(654)

problem is: 2*(15-7*g) = 3*(g+2-3*(2*g-1))
removing left side equation parentheses and right side equation inner parentheses
30 - 14*g = 3*(g + 2 - 6*g + 3)
removing outer parentheses from right side equation
30 - 14*g = 3*g + 6 - 18*g + 9
combining like terms on both sides of equation
30 - 14*g = 15 - 15*g
adding 15*g to both sides of equation and subtracting 30 from both sides of the equation
15*g - 14*g = 15 - 30
combining like terms on both sides of the equation
g = -15
answer is g = -15
-----
original equation
2*(15-7*g) = 3*(g+2-3*(2*g-1)
replacing g with -15 in original equation
2*(15-7*(-15) = 3*((-15)+2 - 3*(2*(-15)-1)
removing parenthese from both sides of equation
30 - 2*7*(-15) = 3*(-15)+3*2 - 3*3*2*(-15)-3*3*(-1)
doing arithmetic
30 - (-210) = -45 + 6 - (-270) - (-9)
240 = -45 + 6 + 270 + 9
240 = 285 - 45
240 = 240
substituting results in original equation is good.
----
answer is g = -15
-----